# Central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discretely valued fields of dimension 2

### v1

@@ Abstract
Let $K$ be a complete discretely valued field with residue field $\bar{K}$ of dimension $1$ (not necessarily perfect). This occurs if and only if $K$ has dimension $2$. We prove the following statements on the arithmetic of such fields:
@@
  • For every prime $p$, every class in the Milnor $\mathrm{K}$-theory modulo $p$ is represented by a symbol.
@@ 1.   Introduction
In this article, we investigate various classical algebraic and arithmetic questions related to central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discrete valuation fields with dimension $2$. Most of the results we provide are well-known when the residue field is assumed to be perfect, and our goal is remove that assumption.
@@
The period-index problem is a classic in the theory of Brauer groups of fields. The period $\mathrm{per}(\alpha)$ of a Brauer class $\alpha$ is its order in the Brauer group and its index $\mathrm{ind}(\alpha)$ is the degree of the division algebra underlying a representative of the class. It is well-known that $\mathrm{per}(\alpha)|\mathrm{ind}(\alpha)$ and that both have the same prime factors (cf. for instance [GS17, Thm. 2.8.7]). Relations of the type $\mathrm{ind}(\alpha)|\mathrm{per}(\alpha)^{n}$ are known for many fields having a nice algebraic, geometric or arithmetic behaviour (e.g. [Art87, AAIKL, Ben24, BHN32, CTOP02, dJ04, FS89, Lie15, Sal97, PS14]). It is often reasonable to expect the equality $\mathrm{ind}(\alpha)=\mathrm{per}(\alpha)$ for Brauer classes $\alpha$ over a field with dimension $2$. In §3, we use an argument due to Kato to prove that such a result holds over any complete discrete valuation field $K$ with dimension $2$ (cf. Propositions 3.1 and 3.2). In the case where $K$ has characteristic 00 and contains all $p$-th roots of unity, where $p$ is the characteristic exponent of its residue field, arguments of Saltman allow us to check a stronger result, stating that all central simple algebras over $K$ are cyclic.
@@
sending the class of the symbol $\{x,y\}$ to the class of the cyclic algebra $(x,y)_{\zeta_{p}}$. This tells us that the cyclicity of central simple algebras of period $p$ can be restated in terms of Milnor $\mathrm{K}$-theory as every class in $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ being represented by a symbol. One could expect then such a result for complete discretely valued fields of dimension 2 without the assumption on roots of unity. One of the main results of this article, proved in §4, corresponds to this statement for fields of characteristic 00:
@@
In order to tackle this statement, the main tool is a description of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic given by Kato. However, this description works under the hypothesis that $K$ contains the $p$-th roots of unity. We remove this assumption in Proposition 4.3. Theorem 1.1 readily follows by analogy with Saltman’s arguments, which we present in this setting in the appendix.
Serre’s Conjecture II
@@
This conjecture remains open, although a lot has been done in particular cases, whether specifying the type of the group (cf. for instance [BP95, BFT07, Gar01, KMRT98, Gar10, Che03, Gil01]) or the field of definition (cf. [Kne65a, Kne65b, BT87, Har65, Har66, Che89, CTGP04, dJHS11, Ben24]. For a survey on the topic, see the lecture notes [Gil19]. More recently, in [ILA24], it is proved that Serre’s Conjecture II can be reduced to the case of countable characteristic 00 fields of dimension $2$.
@@
When the residue field $\bar{K}$ is perfect, this result was already known by work of Kneser in the p-adic case [Kne65a, Kne65b] and Bruhat and Tits [BT87, Cor. 3.15]. The cases of groups of classical type and of types $F_{4}$ and $G_{2}$ were also already known by the aforementioned result of Bayer and Parimala [BP95]. In §6, we deduce the cases of trialitarian $D_{4}$, $E_{6}$ and $E_{7}$ from Propositions 3.1 and 3.2, while we obtain the case of $E_{8}$ by two methods: when $K$ has characteristic 00, we study the maximal solvable extension of $K$ of degree $2^{a}3^{b}5^{c}$ (cf. Proposition 6.2) and use an argument due to the first author; when $K$ has positive characteristic, we follow the original arguments by Bruhat and Tits in [BT87] and adapt them to the case of imperfect residue fields.
@@
In §2, we give some preliminaries on the various existing notions of dimension for fields, on central simple algebras and on Milnor $\mathrm{K}$-theory, recalling in particular an explicit description given by Kato of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic with imperfect residue field. In §3, we settle the “period equals index” property for complete discretely valeud fields of dimension 2 in arbitrary characteristic. In §4, we work in mixed characteristic, we remove the hypothesis on the roots of unity in Kato’s description mentioned before and we apply it in order to prove Theorem 1.1. In order to do so, we state three “absorption lemmas” that help us to reduce sums of symbols to a single symbol (Lemmas 4.4, 4.5 and 4.6). The proof of these lemmas is completely analogous to computations previously carried by Saltman and thus are given in an appendix to this article. In §5, we consider the case where $K$ has positive characteristic and give some partial results in this direction. Finally, in §6, we prove Theorem 1.3 separating the cases of characteristic 00 and positive characteristic.
@@ Acknowledgements
The authors would like to thank Skip Garibaldi for his suggestions that allowed to improve the exposition, as well as Raman Parimala and Suresh Venapally for their interest. The first author was supported by the project “Group schemes, root systems, and related representations” founded by the European Union - NextGenerationEU through Romania’s National Recovery and Resilience Plan (PNRR) call no. PNRR-III-C9-2023-I8, Project CF159/31.07.2023, and coordinated by the Ministry of Research, Innovation and Digitalization (MCID) of Romania. The third author’s research was partially supported by ANID via FONDECYT Grant 1240001.
@@ 2.1 The dimension of a field
The cohomological dimension $\mathrm{cd}(K)$ of a field $K$ is the cohomological dimension of its absolute Galois group. In other words, it is the smallest integer $\delta$ (or $\infty$ if such an integer does not exist) such that $H^{n}(K,M)=0$ for every $n>\delta$ and each finite Galois module $M$. This notion turns out not to be completely satisfactory for imperfect fields, in particular because a field with characteristic $p>0$ and whose Galois group is a pro-$p$-group has cohomological dimension at most $1$. Better definitions in that context were provided by Kato in [Kat82] and by Gille in [Gil00].
@@ 2.3 Milnor $\mathrm{K}$-theory
Let $K$ be any field. The Milnor $\mathrm{K}$-theory groups of $K$ are defined and studied in full generality in [GS17, Ch. 7] (and unless otherwise stated, every statement presented in this subsection comes from there). In this article, we will only use the first three Milnor $\mathrm{K}$-theory groups, which can be explicitly described as:
@@
Milnor $\mathrm{K}$-theory is also endowed with residue maps. Indeed, when $K$ is a complete discretely valued field with ring of integers $\mathcal{O}_{K}$ and residue field $\bar{K}$, there exists a unique residue morphism:
@@
such that, for any uniformizer $\pi$ and for each unit $u\in\mathcal{O}_{K}^{\times}$, one has:
@@ 3.   Period and index of central simple algebras
By [Ser79, XII, §3, Exer. 3] we have an exact sequence:
$0\to{\rm{Br}}(\bar{K})\{\ell\}\to{\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}\to H^{1}(\bar{K},{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})\to 0$ (5)
@@
It remains to deal with the case when $A\otimes_{K}K^{\mathrm{unr}}$ is not split. According to [Ser79, XII, §1, Cor. to Thm. 1], this can happen only in the case where $\bar{K}$ is imperfect and, again by [Ser79, XII, §3, Exer. 3], only if $\ell=p$. Since $\dim_{p}(\bar{K})\leq 1$, we have $[\bar{K}:\bar{K}^{p}]=p$. In this case Kato has shown that $A$ has index $p$ (cf [Kat79, §4, Lemma 5]). ∎
@@ 4.   Symbols and Milnor $K$-theory
The rest of this section is devoted to the proof of this theorem. Let us first deal with the easy case when $\bar{K}$ has characteristic $\neq p$. Under that assumption, since $p$ is invertible in $\bar{K}$, by [GS17, Cor. 7.1.10], we have the exact sequence:
@@
Let us now move on to the case where $\mathrm{char}(\bar{K})=p$. If $K$ contains all $p$-th roots of unity, Theorem 4.1 immediately boils down to Proposition 3.1 since the Bloch-Kato conjecture provides an isomorphism $\mathrm{k}^{\mathrm{M}}_{2}(K)\cong{\rm{Br}}(K)[p]$ in which symbols correspond to cyclic algebras. We henceforth do not assume anymore that $K$ contains all $p$-th roots of unity. In that case, we a priori do not have Kato’s isomorphisms (1)–(4). So we first need to extend them to that case:
@@
$\rho_{i}:H_{p}^{1}(\bar{K})$ $\to u^{i}_{2}(K),$ (9)
$[\bar{x}]$$\mapsto\{1+\pi^{i}ux,\pi\},$
for some well-chosen unit $u\in\mathcal{O}_{K}^{\times}$.
Proof.
@@
When $\zeta_{p}\in K$, the isomorphisms from the statement and the triviality of $u_{2}^{\lfloor\epsilon_{K}\rfloor+1}(K)$ are immediately given by isomorphisms (1)–(4) and the previous computation of $\Omega^{i}_{\bar{K}}$, except that we have reversed the order of the coordinates in some symbols, which amounts to a sign change.
We assume now that $\zeta_{p}\not\in K$ and we define $K^{\prime}:=K(\zeta_{p})$. Then we know that $K^{\prime}/K$ is a totally ramified cyclic extension of degree $d$ dividing $p-1$. More precisely, we know that there exists a uniformizer $\pi\in K$ such that $K^{\prime}=K(\sqrt[d]{\pi})$. Furthermore, we know that isomorphisms (6)-(9) are valid over $K^{\prime}$. We will denote these isomorphisms by $\rho_{i}^{\prime}$.
Write $u_{2}^{0}(K):=\mathrm{k}^{\mathrm{M}}_{2}(K)$ and note that ${\rm{Res\,}}_{K^{\prime}/K}(u_{2}^{i}(K))\subseteq u_{2}^{di}(K^{\prime})$. Thus, there is a natural map
${\rm{Res\,}}_{K^{\prime}/K}:u_{2}^{i}(K)/u_{2}^{i+1}(K)\to u_{2}^{di}(K^{\prime})/u_{2}^{di+1}(K^{\prime}),$ (10)
which we claim to be an isomorphism for every $0\leq i\leq\epsilon_{K}$.
Let us prove surjectivity. This amounts to proving that (the class of) the symbol on the right hand side of $\rho^{\prime}_{di}$ is in the image of the restriction morphism. Now, since these isomorphisms are well-defined regardless of the choice of a uniformizer $\pi^{\prime}$, of the elements $\omega,\varpi\in K^{\prime}$ and of the preimages $x,y\in\mathcal{O}_{K^{\prime}}$, we may choose the uniformizer to be $\pi^{\prime}:=\sqrt[d]{\pi}$ and we may always assume that $x,y\in K$ since $K$ and $K^{\prime}$ have the same residue field.
@@
${\rm{Res\,}}_{K^{\prime}/K}(\{1+\omega x,y\})=\{1+\omega x,y\},$
with $v_{K^{\prime}}(\omega)=di$, so we are done as well.
If $p|i$ with $i>0$ and $p<\epsilon_{K}$, then we have
${\rm{Res\,}}_{K^{\prime}/K}(\{1+\varpi^{p}x,\pi\})=\{1+\varpi^{p}x,{\pi^{\prime}}^{d}\}=d\{1+\varpi_{i}x,\pi^{\prime}\},$
with $v_{K^{\prime}}(\varpi)=di/p$ and we are also done since $(d,p)=1$.
If $i=\epsilon_{K}$, then $(\zeta_{p}-1)^{p}=\pi^{\epsilon_{K}}u_{0}$ for some $u_{0}\in\mathcal{O}_{K^{\prime}}^{*}$. Since $K^{\prime}/K$ is totally ramified, we may fix $u\in\mathcal{O}_{K}^{*}$ such that $\bar{u}_{0}=\bar{u}$. Then,
${\rm{Res\,}}_{K^{\prime}/K}(\{1-\pi^{\epsilon_{K}}ux,\pi\})=\{1-(\zeta_{p}-1)^{p}u_{0}^{-1}ux,{\pi^{\prime}}^{d}\}=d\{1-(\zeta_{p}-1)^{p}x,\pi^{\prime}\},$ (11)

### v2

# Central simple algebras, Milnor $\mathrm{K}$-theory and homogeneous spaces over complete discretely valued fields of dimension 2

@@ Abstract
Let $K$ be a complete discretely valued field of dimension $2$. We prove the following statements on the arithmetic of $K$:
@@
  • If $K$ has characteristic 00, for every prime $p$, every class in the Milnor $\mathrm{K}$-theory modulo $p$ is represented by a symbol.
@@ 1.   Introduction
In this article, we investigate various classical algebraic and arithmetic questions related to central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discrete valuation fields with dimension $2$. Most of the results we provide are well-known when the residue field is assumed to be perfect, and our goal is to remove that assumption. This fills an important gap in the literature, since complete discrete valuation fields with imperfect residue fields are common in arithmetic settings, even when one is interested in studying characteristic zero fields. They arise for instance as completions of finitely generated fields.
@@
The period-index problem is a classic in the theory of central simple algebras and the Brauer groups of fields. The period $\mathrm{per}(\alpha)$ of a Brauer class $\alpha$ is its order in the Brauer group and its index $\mathrm{ind}(\alpha)$ is the degree of the division algebra underlying a representative of the class. It is well-known that $\mathrm{per}(\alpha)|\mathrm{ind}(\alpha)$ and that both have the same prime factors (cf. for instance [GS17, Thm. 2.8.7]). Relations of the type $\mathrm{ind}(\alpha)|\mathrm{per}(\alpha)^{n}$ are known for many fields having a nice algebraic, geometric or arithmetic behaviour (e.g. [Art87, AAIKL, Ben24, BHN32, CTOP02, dJ04, FS89, Lie15, Sal97, PS14]). It is often reasonable to expect the equality $\mathrm{ind}(\alpha)=\mathrm{per}(\alpha)$ for Brauer classes $\alpha$ over a field with dimension $2$. From this perspective and thanks to an argument of Kato, we prove that such a result holds over any complete discrete valuation field $K$ with dimension $2$ (cf. Propositions 3.1 and 3.2). Moreover, in the case where $K$ has characteristic 00 and contains all $p$-th roots of unity, where $p$ is the characteristic exponent of its residue field, arguments of Saltman allow us to check a stronger result, stating that all central simple algebras over $K$ are cyclic.
@@
sending the class of the symbol $\{x,y\}$ to the class of the cyclic algebra $(x,y)_{\zeta_{p}}$. In particular, the cyclicity of central simple algebras of period $p$ over $K$ can be restated in terms of Milnor $\mathrm{K}$-theory as every class in $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ being represented by a symbol. It is then natural to ask whether such a result holds for complete discretely valued fields of dimension 2 without the assumption on roots of unity. The first main result of this article, proved in §4, shows this for fields of characteristic 00:
@@
In order to tackle this statement, the main tool is a description of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic given by Kato. However, this description works under the hypothesis that $K$ contains the $p$-th roots of unity. We remove this assumption in Proposition 4.3. Theorem 1.1 then readily follows by analogy with Saltman’s arguments, which we present in this setting in the appendix.
Homogeneous spaces.
@@
This conjecture remains open, although a lot has been done in particular cases, whether specifying the type of the group (cf. for instance [BP95, BFT07, Gar01, KMRT98, Gar10, Che03, Gil01]) or the field of definition (cf. [Kne65a, Kne65b, BT87, Har65, Har66, Che89, CTGP04, dJHS11, Ben24]). For a survey on the topic, see the lecture notes [Gil19]. More recently, in [ILA24], it is proved that Serre’s Conjecture II can be reduced to the case of countable characteristic 00 fields of dimension $2$.
@@
When the residue field $\bar{K}$ is perfect, this result was already known by work of Kneser in the $p$-adic case [Kne65a, Kne65b] and Bruhat and Tits [BT87, Cor. 3.15]. The cases of groups of classical type and of types $F_{4}$ and $G_{2}$ were also already known by a result of Bayer and Parimala [BP95]. In §6, we deduce the cases of trialitarian $D_{4}$, $E_{6}$ and $E_{7}$ from Propositions 3.1 and 3.2, while we obtain the case of $E_{8}$ by two methods: when $K$ has characteristic 00, we study the maximal solvable extension of $K$ of degree $2^{a}3^{b}5^{c}$ (cf. Proposition 6.2) and use an argument due to the first author; when $K$ has positive characteristic, we follow the original arguments by Bruhat and Tits in [BT87] and adapt them to the case of imperfect residue fields.
This result has several applications. The first one we should mention is that Theorem 6.1 permits to extend to the $E_{8}$-case the method of Kisin-Pappas on extension of torsors [KP18, §1.4] and consequently several other results, see §7. We also give applications to the study of torsors over curves in Corollaries 6.5 and 6.6. Finally, in the characteristic zero case, we generalize Kneser’s computation of the Galois cohomology of an arbitrary semisimple group, see Theorem 6.7.
@@
In §2, we give some preliminaries on the various existing notions of dimension for fields, on central simple algebras and on Milnor $\mathrm{K}$-theory, recalling in particular an explicit description given by Kato of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic with imperfect residue field. In §3, we settle the “period equals index” property for complete discretely valeud fields of dimension 2 in arbitrary characteristic. In §4, we work in mixed characteristic, we remove the hypothesis on the roots of unity in Kato’s description mentioned before and we apply it in order to prove Theorem 1.1. In order to do so, we state three “absorption lemmas” that help us to reduce sums of symbols to a single symbol (Lemmas 4.5, 4.6 and 4.7). The proof of these lemmas is completely analogous to computations previously carried by Saltman and thus are given in an appendix to this article. In §5, we consider the case where $K$ has positive characteristic and give some partial results in this direction. Finally, in §6, we prove Theorem 1.3 separating the cases of characteristic 00 and positive characteristic. We immediately deduce several applications, which are scattered through §6 and §7.
@@ Acknowledgements
The authors would like to thank Skip Garibaldi for his suggestions that allowed to improve the exposition, as well as Georgios Pappas, Raman Parimala and Suresh Venapally for their interest.
The first author was supported by the project “Group schemes, root systems, and related representations” founded by the European Union - NextGenerationEU through Romania’s National Recovery and Resilience Plan (PNRR) call no. PNRR-III-C9-2023-I8, Project CF159/31.07.2023, and coordinated by the Ministry of Research, Innovation and Digitalization (MCID) of Romania. The third author’s research was partially supported by ANID via FONDECYT Grant 1240001.
@@ 2.1 The dimension of a field
The cohomological dimension $\mathrm{cd}(K)$ of a field $K$ is the cohomological dimension of its absolute Galois group. In other words, it is the smallest integer $\delta$ (or $\infty$ if such an integer does not exist) such that $H^{n}(K,M)=0$ for every $n>\delta$ and each finite Galois module $M$. This notion turns out not to be completely satisfactory for imperfect fields, in particular because a field with characteristic $p>0$ and whose Galois group is a pro-$p$-group has cohomological dimension at most $1$. Better definitions in that context were provided by Kato in [Kat82] and by the first author in [Gil00].
@@ 2.3 Milnor $\mathrm{K}$-theory
Let $K$ be any field. The Milnor $\mathrm{K}$-theory groups of $K$ are defined and studied in full generality in [GS17, Ch. 7] (and unless otherwise stated, every statement presented in this subsection comes from there). In this article, we will only use the first three Milnor $\mathrm{K}$-theory groups, which can be explicitly described as follows:
@@
Milnor $\mathrm{K}$-theory is also endowed with residue maps. Indeed, when $K$ is a complete discretely valued field with ring of integers $\mathcal{O}_{K}$ and residue field $\bar{K}$, there exists a unique residue morphism
@@
such that, for any uniformizer $\pi$ and for each unit $u\in\mathcal{O}_{K}^{\times}$, one has
@@ 3.   Period and index of central simple algebras
By [Ser79, XII, §3, Exer. 3] we have an exact sequence
$0\to{\rm{Br}}(\bar{K})\{\ell\}\to{\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}\to H^{1}(\bar{K},{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})\to 0$
@@
It remains to deal with the case when $A\otimes_{K}K^{\mathrm{unr}}$ is not split. According to [Ser79, XII, §1, Cor. to Thm. 1], this can happen only in the case where $\bar{K}$ is imperfect and, again by [Ser79, XII, §3, Exer. 3], only if $\ell=p$. Since $\dim_{p}(\bar{K})\leq 1$, we have $[\bar{K}:\bar{K}^{p}]=p$. In this case Kato has shown that $A$ has index $p$ (cf. [Kat79, §4, Lemma 5]). ∎
@@ 4.   Symbols and Milnor $K$-theory
The rest of this section is devoted to the proof of this theorem. Let us first deal with the easy case when $\bar{K}$ has characteristic $\neq p$. Under that assumption, since $p$ is invertible in $\bar{K}$, by [GS17, Cor. 7.1.10], we have the exact sequence
@@
Let us now move on to the case where $\mathrm{char}(\bar{K})=p$. If $K$ contains all $p$-th roots of unity, Theorem 4.1 immediately boils down to Proposition 3.1 since the Bloch-Kato conjecture provides an isomorphism $\mathrm{k}^{\mathrm{M}}_{2}(K)\cong{\rm{Br}}(K)[p]$ in which symbols correspond to cyclic algebras. We henceforth do not assume anymore that $K$ contains all $p$-th roots of unity. In that case, we a priori do not have Kato’s isomorphisms (1)–(4). So we first need to extend them to that case.
@@
$\rho_{i}:\bar{K}/(\bar{K}\cap\overline{u}\Phi(\bar{M}))$ $\to u^{i}_{2}(K),$ (8)
$[\bar{x}]$$\mapsto\{1+\omega x,\pi\}$
where $\omega$ is an element in $K$ with $v_{K}(\omega)=i$, the field $M$ stands for $K(\zeta_{p})$, $u$ is an element of $\mathcal{O}_{M}^{\times}$ such that $\omega u=(\zeta_{p}-1)^{p}$ and $\Phi(\bar{M}):=\{t^{p}-t:t\in\bar{M}\}$.
Remark 4.4.
When $\zeta_{p}\in K$, i.e. when $M=K$, the last isomorphism recovers Kato’s isomorphism when taking $\omega=(\zeta_{p}-1)^{p}$. Indeed, then $\bar{u}=1$ and one checks that the quotient $\bar{K}/(\bar{K}\cap\Phi(\bar{K}))$ is precisely $H^{1}_{p}(\bar{K})$.
More generally, when $M/K$ is totally ramified, we have $\bar{u}\in\bar{M}=\bar{K}$, hence the map
$H^{1}_{p}(\bar{K})=\bar{K}/\Phi(\bar{K})\rightarrow\bar{K}/(\bar{K}\cap\bar{u}\Phi(\bar{K})),\bar{x}\mapsto\bar{u}\bar{x}$
is an isomorphism and $\rho_{\epsilon_{K}}$ induces an isomorphism between $H^{1}_{p}(\bar{K})$ and $u_{2}^{i}(K)$.
On the other hand, when $M/K$ is unramified, we may have that $\bar{K}/(\bar{K}\cap\overline{u}\Phi(\bar{M}))$ is non isomorphic to $H^{1}_{p}(\bar{K})$, even when $\bar{K}$ is perfect. For instance, if $K={\mathbb{Q}}_{3}(\sqrt{-6})$, then $\bar{K}={\mathbb{F}}_{3}$ and $\bar{M}={\mathbb{F}}_{9}$. A direct computation gives that
$H_{p}^{1}(\bar{K})=\bar{K}/\Phi(\bar{K})={\mathbb{F}}_{3}.$
Now, if we choose $\pi=\sqrt{-6}$ as a uniformizer and $\omega:=\pi^{3}$, we get that $\bar{u}=\sqrt{2}^{-1}\in\bar{M}$. Since $\Phi(\sqrt{2})=\sqrt{2}$, we get $1\in\bar{u}\Phi(\bar{M})$ and hence
$\bar{K}/(\bar{K}\cap\bar{u}\Phi(\bar{M}))=\bar{K}/\bar{K}=0.$
Proof of Proposition 4.3.
@@
When $\zeta_{p}\in K$, the isomorphisms from the statement and the triviality of $u_{2}^{\lfloor\epsilon_{K}\rfloor+1}(K)$ are immediately given by isomorphisms (1)–(4), Remark 4.4 and the previous computation of $\Omega^{i}_{\bar{K}}$, except that we have reversed the order of the coordinates in some symbols, which amounts to a sign change.
We assume now that $\zeta_{p}\not\in K$. Define $M:=K(\zeta_{p})$ and $L/K$ is the maximal unramified extension of $M/K$. Then we know that $M/L$ is a totally ramified cyclic extension of degree $d$ dividing $p-1$. More precisely, we know that there exists a uniformizer $\pi\in K$ such that $M=L(\sqrt[d]{\pi})$. Furthermore, we know that isomorphisms (5)-(8) are valid over $M$. We will denote these isomorphisms by $\rho_{i}^{M}$.
Let us assume first that $L=K$. Write $u_{2}^{0}(K):=\mathrm{k}^{\mathrm{M}}_{2}(K)$ and note that, for every $i\geq 0$, ${\rm{Res\,}}_{M/K}(u_{2}^{i}(K))\subseteq u_{2}^{di}(M)$. Thus, there is a natural map
${\rm{Res\,}}_{M/K}:u_{2}^{i}(K)/u_{2}^{i+1}(K)\to u_{2}^{di}(M)/u_{2}^{di+1}(M),$ (9)
which we claim to be an isomorphism for every $0\leq i\leq\epsilon_{K}$. Moreover, up to multiplication by an integer prime to $p$, its inverse is given by the norm map $N_{M/K}$.
Let us prove surjectivity. This amounts to proving that (the class of) the symbol on the right hand side of $\rho^{M}_{di}$ is in the image of the restriction morphism. Now, since these isomorphisms are well-defined regardless of the choice of a uniformizer $\pi_{M}$, of the elements $\omega,\varpi\in M$ and of the preimages $x,y\in\mathcal{O}_{M}$, we may choose the uniformizer to be $\pi_{M}:=\sqrt[d]{\pi}$ and we may always assume that $x,y\in K$ since $K$ and $M$ have the same residue field.
@@
${\rm{Res\,}}_{M/K}(\{1+\omega x,y\})=\{1+\omega x,y\},$
with $v_{M}(\omega)=di$, so we are done as well.
If $p|i$ with $i>0$ and $i<\epsilon_{K}$, then we have
${\rm{Res\,}}_{M/K}(\{1+\varpi^{p}x,\pi\})=\{1+\varpi^{p}x,\pi_{M}^{d}\}=d\{1+\varpi^{p}x,\pi_{M}\},$
with $v_{M}(\varpi)=di/p$ and we are also done since $(d,p)=1$.
If $i=\epsilon_{K}$, then $(\zeta_{p}-1)^{p}=\pi^{\epsilon_{K}}u_{0}$ for some $u_{0}\in\mathcal{O}_{M}^{\times}$. Since $M/K$ is totally ramified, we may fix $u\in\mathcal{O}_{K}^{\times}$ such that $\bar{u}_{0}=\bar{u}$. Then,
${\rm{Res\,}}_{M/K}(\{1-\pi^{\epsilon_{K}}ux,\pi\})=\{1-(\zeta_{p}-1)^{p}u_{0}^{-1}ux,\pi_{M}^{d}\}=d\{1-(\zeta_{p}-1)^{p}x,\pi_{M}\},$